OpenStudy (anonymous):

A horse pulls on an object with a force of 300 newtons and does 12000 joules of work. How far was the object moved ?

OpenStudy (michele_laino):

we have to apply this formula: \[W = Fd\] where W is work done by the horse, F is the force exerted by the horse, and d is the distance

OpenStudy (michele_laino):

so, we have: \[d = \frac{W}{F}\]

OpenStudy (michele_laino):

\[d = \frac{W}{F} = \frac{{12000}}{{300}} = ...Newtons\]

OpenStudy (anonymous):

word equals force times distance

OpenStudy (pawanyadav):

Distance=work/force